Apply Charpit's method to solve the equation z-px-qy=p^2+q^2
1
Expert's answer
2020-12-16T14:51:07-0500

Given differential equation is "z-px-qy-p^2-q^2=0"

Then, "f(x,y,z,p,q) = z-px-qy-p^2-q^2=0"


"\frac{dx}{-\frac{\partial f}{\partial p}} = \frac{dy}{-\frac{\partial f}{\partial q}}=\frac{dz}{-p\frac{\partial f}{\partial p}-q\frac{\partial f}{\partial q}}=\frac{dp}{\frac{\partial f}{\partial x}+p\frac{\partial f}{\partial z}}=\frac{dq}{\frac{\partial f}{\partial y}+q\frac{\partial f}{\partial z}}"


Then,

"\frac{\partial f}{\partial p} = -(x+2p)", "\frac{\partial f}{\partial q}=-(y+2q)", "\frac{\partial f}{\partial x} =-p" , "\frac{\partial f}{\partial y}=-q" , "\frac{\partial f}{\partial z} = 1"


Putting values, we get,

"\frac{dx}{x+2p} = \frac{dy}{y+2q}=\frac{dz}{p(x+2p)+q(y+2q)}=\frac{dp}{-p+p}=\frac{dq}{-q+q}"



"\frac{dx}{x+2p} = \frac{dy}{y+2q}=\frac{dz}{px+qy+2p^2+2q^2}=\frac{dp}{0}=\frac{dq}{0}"


Taking second last term, "dp = 0 \implies p = a" (1)


taking last term, "dq=0 \implies q=b" (2)


Then, "dz = pdx+qdy = adx+bdy"

integrating both sides, "z = ax+by+c" (3)


Now, equation will be,

"ax+by+c = ax+by+a^2+b^2"


"\implies c = a^2+b^2"


Hence, "z = ax+by+a^2+b^2"